If RS=23-2x,ST=9x-5,and RT=39 find RS.
RS=

If LN=6x-5,LM=x+7, and MN=3x+20 find MN
MN

Find X Round tithe nearest hundredth.
X(-9,2) and Y(5,-4)

Find the midpoint of AB if A(-3,8) and B(-7,-6)

If CE=10x+18,DE=7x-1 and BC=9x-2, find AB
X=
AB=
1
Expert's answer
2020-09-22T15:53:27-0400

1.If RS=23-2x,ST=9x-5,and RT=39 find RS.

"rt = rs + st"

rt = 39

rs = 23 – 2x

st = 9x-5

"39 = (23 \u2013 2x) + (9x - 5)"

"x = 3"

"rs = 23 - 2 * 3"

RS = 17


2.If LN=6x-5,LM=x+7, and MN=3x+20 find MN

"6*x - 5 = x + 7 + 3*x + 20"

"6x - 4x = 32"

"x = 16"

"MN = 3*16+20"

MN = 68


3.Find X Round tithe nearest hundredth.

X(-9,2) and Y(5,-4)

"d = \sqrt{ (5-(-9))^2 + (-4-2)^2}"

answer: 15.23


4.Find the midpoint of AB if A(-3,8) and B(-7,-6)

"midpoint = ((-3-7)\/2; (8-6)\/2)"

answer: (-5 , 1)


5.If CE=10x+18, DE=7x-1 and BC=9x-2, find AB

"CE=10x+18"

"DE=7x-1"

D is midpoint

"DE = 1\/2CE"

"7*x - 1 = 1\/2*(10*x+18)"

X=5

"CE = 10*5+18 = 68 = AC"

"BC = 9*x-2 = 9*5-2 = 43"

AB = AC - BC = 68-43 = 25

AB=25